OpenStudy (anonymous):

@TheSmartOne An activity takes 63 minutes when there are 4 participants. The activity takes 87 minutes when there are 6 participants. Assume the relationship is linear. A line models the relationship between the number of participants, x, and the number of minutes, y. What is the slope of the line?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Woops I posted the wrong question. Hold on...

OpenStudy (anonymous):

OK, I changed the question, do you know this one? :)

OpenStudy (aum):

You are given two points on a line: (4,63) and (6,87). Find the slope. Then find the equation of a line through it.

OpenStudy (aum):

Oh, they just want the slope. So you can stop there.

OpenStudy (aum):

\[ \text{Slope } = \frac{y_2-y_1}{x_2-x_1} \]

OpenStudy (anonymous):

So...\[\frac{ 4-6 }{ 63-87 } =\frac{ -2 }{ -24 } = ?\]

OpenStudy (anonymous):

@aum

OpenStudy (aum):

You have reversed the x and y's. It should be the reciprocal.

OpenStudy (aum):

\[ \frac{87-63}{6-4} = ? \]

OpenStudy (anonymous):

Oh woops. :) \[\frac{ 24 }{ 2 } = 12?\]

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Yes, slope = 12.